OpenStudy (anonymous):

$1000 loan, 6% interest rate, for 11/2yr, compound semi-annually, how much interest did pay?

OpenStudy (anonymous):

\[I = P(1 + \frac rn)^{nt}\]

OpenStudy (anonymous):

P=1000(1+0.03)^3 P=1092.727 interest equals P-1000= $92.727

OpenStudy (anonymous):

\[1000*(1+[0.06^{11/2}\])-1000\

OpenStudy (anonymous):

Thank you all for your help!!

OpenStudy (anonymous):

Could you tell me why wee would add 1 + 0.03?

OpenStudy (anonymous):

the 0.03 is the rate of the half year compounded rate adn one just is just the repeat of the 1000 dollars

OpenStudy (anonymous):

o im wrong for semi-annually interest rate is usually for a year

OpenStudy (anonymous):

That's alright AndrewNJ, thanks for your help anyway

OpenStudy (anonymous):

Thank you Calcmathlete

OpenStudy (anonymous):

Thank you aceace

OpenStudy (anonymous):

You're welcome :)